A particle with a charge of 38 μC moves with a speed of 73 m/s in the positive x direction. The magnetic field in this region of space has a component of 0.44 T in the positive y direction, and a component of 0.84 T in the positive z direction.

What is the magnitude of the magnetic force on the particle?

What is the direction of the magnetic force on the particle? (Find the angle measured from the positive z-axis toward the negative y-axis in the yz-plane.)
Answer #1

V=73 mis . Ba 0.44 T Ĵ + 0.ght @ E = q (NXB) = 38*4** * = 38 xioo (73ix (0.44f + 0.84 it) 38x18t 192.12 Å - 661 32 ŷ ) 2 -22